OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. 2x – y + z = –3 2x + 2y + 3z = 2 3x – 3y – z = –4 (–1, –3, –2) (1, –3, –2) (1, 3, –2) (1, 3, 2)

OpenStudy (anonymous):

Combine the first and third equations and eliminate z variables: 2x – y + z = –3 3x – 3y – z = –4 5x - 4y = -7 Combine the first two equations and also eliminate z variables by multiplication: -3(2x – y + z = –3) 2x + 2y + 3z = 2 -6x + 3y - 3z = 9 2x + 2y + 3z = 2 -4x + 5y = 11 Here, we have: -4x + 5y = 11 5x - 4y = -7 Then, 4(-4x + 5y = 11) 5(5x - 4y = -7) -16x + 20y = 44 25x - 20y = -35 9x = 9 x = 1 Substitute that value for either two-variable equation and solve for y: 5(1) - 4y = -7 5 - 4y = -7 -4y = -12 y = 3 Finally, substitute x and y values for either three-variable equation and solve for z! 2(1) – 3 + z = –3 2 - 3 + z = -3 -1 + z = -3 z = -2 Hence, the solution is (1,3,-2)

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

No problem :)